# Visual Perturbation and Adaptive Hard Negative Contrastive Learning for Compositional Reasoning in Vision-Language Models

### v1

@@ 4.6 Visualization
To demonstrate the clear effectiveness of our proposed visual perturbation strategy, a compelling visualization study is provided. The Fig. 5 shows two illustrative examples, where each example generates textual negative samples using the methods in Section 3.1. Specifically, for each example, we extract $e_{I_{orig}}$, $e_{T_{orig}}$ and $e_{T_{neg}}$ using zero-shot CLIP, AHNPL at epoch 5 and 10 (training end), respectively. The visual negative embedding $e_{I_{neg}}$ is then calculated from these extracted embeddings via our proposed visual perturbation. We illustrate the cosine distances between key pairs of these four embeddings. The changes in cosine distances demonstrate that the visual perturbation strategy effectively pushes the original image features towards negative texts, improving feature distinguishability and thereby helping the model capture subtle positive-negative semantic differences.
Figure 5: Impact of AHNPL training on the embedding space distances of two illustrative examples. Numerical values represent cosine distances between key embedding pairs.
@@ 5 Conclusion
This paper has proposed AHNPL, enhancing VLMs performance in CR tasks via Visual Perturbation with Semantic Shift and Dynamic Hard Negative Contrastive Learning. AHNPL transforms text-based hard negatives into semantically perturbed image negatives, effectively improving the overall performance. Meanwhile, Dynamic Hard Negative Contrastive Learning strengthens the model’s ability to align positive samples and distinguish hard negatives. Experimental results show that AHNPL outperforms existing methods in downstream tasks, showing robust complex semantic understanding in multimodal scenarios. Future work will explore integrating large-scale knowledge graphs and implicit knowledge to enhance complex semantic relationship acquisition.

### v2

@@ 4.6 Visualization
To demonstrate the clear effectiveness of our proposed visual perturbation strategy, a compelling visualization study is provided. Tab. 5 shows two illustrative examples, where each example generates textual negative samples using the methods in Section 3.1. Specifically, for each example, we extract $e_{I_{orig}}$, $e_{T_{orig}}$ and $e_{T_{neg}}$ using zero-shot CLIP, AHNPL at epoch 5 and 10 (training end), respectively. The visual negative embedding $e_{I_{neg}}$ is then computed from these extracted embeddings via our proposed visual perturbation. We illustrate the cosine distances between key pairs of these four embeddings. The changes in cosine distances demonstrate that the visual perturbation strategy effectively pushes the original image features towards negative texts, improving feature distinguishability and thereby helping the model capture subtle positive-negative semantic differences.
Table 5: Impact of AHNPL training on the embedding space distances of two illustrative examples. Numerical values represent cosine distances between key embedding pairs.
@@ 5 Conclusion
This paper proposes AHNPL, enhancing VLMs’ performance in CR tasks via Visual Perturbation with Semantic Shift and Dynamic Hard Negative Contrastive Learning. AHNPL transforms text-based hard negatives into semantically perturbed image negatives, effectively improving the overall performance. Additionally, Dynamic Hard Negative Contrastive Learning strengthens the model’s ability to align positive samples and distinguish hard negatives. Experimental results show that AHNPL outperforms existing methods in downstream tasks, showing robust complex semantic understanding in multimodal scenarios. Future work will explore integrating large-scale knowledge graphs and implicit knowledge to enhance complex semantic relationship acquisition.